OpenStudy (anonymous):

solve for x, y and/or z (solution to 2 linear equations) \[\left(\begin{matrix}3x+2y=4 \\ 2x-4y=1\end{matrix}\right)\]

OpenStudy (anonymous):

Does it matter if you use substition or elemination?

OpenStudy (anonymous):

i think elimination is the method to be used

OpenStudy (anonymous):

(3x + 2y = 4)(2) 2x - 4y = 1 ---------------------- 5x = 9 x = 5/9 y = 6/5

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

your welcome :-)